### v2

@@ Discussion
To gain physical insight into the classification process, we reconstructed the LDA discriminant vectors into image space. These projection vectors, shown in Fig. 6(c), highlight spatial and spectral regions most relevant for discrimination. For example, in the 30D PCA case of Fig. 6(c), the green channel is stronger at the image center, while the red channel is weaker. Combined with the kernel density estimation (KDE) plots in Fig. 6(d)—where spheres have higher projection scores than cubes—this indicates that, statistically, nanospheres exhibit higher green intensity at the center of their DF images, whereas nanocubes display more pronounced red in the same region. Physically, when the overall dimensions are similar, this difference may arise from the distinct far-field distributions of the two geometries: nanospheres tend to support more symmetric scattering modes that enhance shorter-wavelength (green) scattering at the center, while nanocubes, with sharper edges and corners, can localize fields in a way that favors longer-wavelength (red) scattering in the central region [59]. These projection vectors can also be used directly for classification by correlating them with normalized DF images and applying a single decision threshold (Fig. 6(d)), effectively combining feature extraction and classification in one interpretable step.
@@
However, our DL-based SPARX reduces both cost and technical barriers: after training, it relies solely on standard dark-field microscopy components (e.g., an RGB camera) and does not require a spectrometer, whereas conventional hyperspectral systems typically demand more expensive, specialized hardware and stricter spectral–spatial calibration. This positions SPARX as an accessible, lab-ready solution for general optical characterization and specialized plasmonic applications—validated across diverse nanoplasmonic geometries, including nanospheres and nanocubes as demonstrated in this work. And SPARX is compatible with emerging techniques such as compressed sensing, enabling further acceleration when dealing with sparse samples.
@@ Methods
To analyze our data, we employ PCA, a statistical technique that reduces high-dimensional data into a lower-dimensional space by identifying new orthogonal axes—called principal components—that capture the maximum variance within the dataset. This transformation helps reveal patterns and correlations that may not be obvious in the original feature space. Since DF images are in RGB format, the PCA components of the DF images are normalized between 0 and 1 for better color visualization. To assess the reliability and structure of the DF image PCA projections, we applied several color mappings (see Figure 2):
@@
    Spectral features: We also used two key spectral metrics for coloring—(i) the resonance wavelength ($\lambda$), and (ii) the total spectral energy—both of which show that clusters in the DF projection map correspond to physically meaningful spectral properties.
@@
Electromagnetic simulations were done with the commercial finite-element method package COMSOL Multiphysics. Since the nanoparticle on mirror system has an axis-symmetry, we implemented so-called 2.5D calculation method [5] on that with an oblique incident light (considering 0.9 NA). Permittivity of the gold follows the famous Olmon et.al, dataset [66]. The scattering spectra of the nanoparticles were calculated by integrating the energy flow of the scattered field. We considered that the molecule layer has a refractive index of 1.4 situated between the nanoparticle and the film to acting as insulator to form a metal-insulator-metal nanocavity.
@@
[…]
Acknowledgments This work was supported by the National Key Research and Development Program of China (Grant No. 2024YFA1409900) and the National Natural Science Foundation of China (Grant No. 62475071 and 52488301). M.K., and F.P. acknowledge funding by European Union’s Horizon Europe under grant agreement 101125498, project MINING - “Multifunctional nano-bio interfaces with deep brain regions”. M.K., and F. P. acknowledge funding from the European Union’s Horizon 2020 Research and Innovation Programme project DEEPER under grant agreement 101016787.
[…]
Conflict of interests The authors declare no competing interests.
[…]
Authors’ contributions M. Kazemzadeh and B. Zhang contributed equally to this work. H. Hu conceived the idea and did electromagnetic simulations. M. Kazemzadeh trained the deep learning model and analyzed the data. B. Zhang, W. Jiang, and W. Chen performed the experimental investigation and collected data for the training. W. Chen, B. Zhang, T. He, H. Liu, Z. Jiang, Z. Hu, X. Dong, C. Sun, X. He, and H. Xu contributed to the establishment of the automatic collection optical setup. M. Kazemzadeh and H. Hu created the figures and wrote the original draft, reviewed and edited. All the authors discussed, reviewed, and edited the paper.
[…]
Data availability Source data, codes, and models are available at https://doi.org. All other data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
[…]
Ethics approval Not applicable.

### v3

@@ Introduction
To gain physical insight into the classification process, we reconstructed the LDA discriminant vectors into image space. These projection vectors, shown in Fig. 6(c), highlight spatial and spectral regions most relevant for discrimination. For example, in the 30D PCA case of Fig. 6(c), the green channel is stronger at the image center, while the red channel is weaker. Combined with the kernel density estimation (KDE) plots in Fig. 6(d) where spheres have higher projection scores than cubes. This indicates that, statistically, nanospheres exhibit higher green intensity at the center of their DF images, whereas nanocubes display more pronounced red in the same region. Physically, when the overall dimensions are similar, this difference may arise from the distinct far-field distributions of the two geometries: nanospheres tend to support more symmetric scattering modes that enhance shorter-wavelength (green) scattering at the center, while nanocubes, with sharper edges and corners, can localize fields in a way that favors longer-wavelength (red) scattering in the central region 13. These projection vectors can also be used directly for classification by correlating them with normalized DF images and applying a single decision threshold (Fig. 6(d)), effectively combining feature extraction and classification in one interpretable step.
@@
However, our DL-based SPARX reduces both cost and technical barriers: after training, it relies solely on standard dark-field microscopy components (e.g., an RGB camera) and does not require a spectrometer, whereas conventional hyperspectral systems typically demand more expensive, specialized hardware and stricter spectral–spatial calibration. While variations in microscope optics, illumination, and camera response can affect measured RGB values, a limitation relevant for generalizability and commercial deployment, this issue can be largely mitigated in controlled laboratory setups using standardized optical components, which are common in commercial systems. Additional strategies such as transfer learning with small calibration datasets or domain-adaptive normalization could further enhance cross-hardware robustness. Furthermore, SPARX leverages spatial intensity distributions and inter-channel correlations, which are inherently more resilient to hardware variations than absolute RGB values alone. This positions SPARX as an accessible, lab-ready solution for general optical characterization and specialized plasmonic applications—validated across diverse nanoplasmonic geometries, including nanospheres and nanocubes as demonstrated in this work. And SPARX is compatible with emerging techniques such as compressed sensing, enabling further acceleration when dealing with sparse samples.
@@ Methods
Supervised Learning
We develop a hybrid 2D–1D deep neural network to predict one-dimensional spectral responses from two-dimensional dark-field scattering images. The model takes $128\times 128\times 3$ images as input and outputs a spectral sequence of length 128, along with associated uncertainty estimates.The network follows an encoder–decoder architecture. The encoder consists of six hierarchical 2D convolutional stages, each comprising a residual convolutional block and a max-pooling layer (stride $2\times 2$) for progressive multi-scale feature extraction. Each residual block is built from stacked $3\times 3$ convolutions with Batch Normalization and ReLU activation, and incorporates L1–L2 regularization ($l_{1}=10^{-5},\;l_{2}=10^{-4}$) to enhance generalization.At the end of the encoder, high-dimensional features are flattened and projected through a fully connected layer, then reshaped into a one-dimensional latent representation, enabling the transformation from spatial to spectral domain.
The decoder comprises five 1D convolutional stages. Each stage applies a residual Conv1D block (kernel size 3) followed by transposed convolution (stride 2) for progressive upsampling, recovering the output sequence to length 128. A final Conv1D layer produces a two-channel output corresponding to the predicted spectrum and its uncertainty.
The dataset was divided into 8,382 images for training, 2,794 images for validation during training, and an independent test set of 1,000 images with paired spectral measurements used exclusively for final evaluation. The model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$, a batch size of 32, and 150 epochs. An adaptive learning rate scheduling strategy based on validation performance is employed to facilitate convergence.
To mitigate over-fitting, model training employed validation-based checkpointing and early stopping, with the best-performing model selected according to the minimum validation loss. Consistent training and validation losses and stable performance on an independent test set indicate good generalization.
For the classification study, nanoparticle labels were assigned according to the fabrication method and the shape and size distributions provided by the commercial supplier. Minor polydispersity or irregular by-products may be present but were not considered in this classification study.
Please find the source code, trained models, and the curated dataset available from the repository stated in the Data availability section.
@@
To analyze our data, we employ PCA, a statistical technique that reduces high-dimensional data into a lower-dimensional space by identifying new orthogonal axes, principal components, that capture the maximum variance within the dataset. This transformation helps reveal patterns and correlations that may not be obvious in the original feature space. Since DF images are in RGB format, the PCA components of the DF images are normalized between 0 and 1 for better color visualization. To assess the reliability and structure of the DF image PCA projections, we applied several color mappings (see Figure 2):
@@
    Spectral features: We also used two key spectral metrics for coloring (i) the resonance wavelength ($\lambda$), and (ii) the total spectral energy. Both of which show that clusters in the DF projection map correspond to physically meaningful spectral properties.
@@
Electromagnetic simulations were done with the commercial finite-element method package COMSOL Multiphysics. Since the nanoparticle on mirror system has an axis-symmetry, we implemented so-called 2.5D calculation method 14 on that with an oblique incident light (considering 0.9 NA). Permittivity of the gold follows the famous Olmon et.al, dataset 47. The scattering spectra of the nanoparticles were calculated by integrating the energy flow of the scattered field. We considered that the molecule layer has a refractive index of 1.4 situated between the nanoparticle and the film to acting as insulator to form a metal-insulator-metal nanocavity.
@@
Gold mirror was fabricated using the template-stripped method 45. Briefly, 100-nm-thick Au films were thermally evaporated onto a clean silicon wafers ($\sim{0.5}\;\rm{nm\cdot s^{-1}}$ deposition rate) followed by epoxy bonding to quartz substrates using UV-curable optical adhesive (NOA61, Norland Products). Mechanical cleavage at the silicon-quartz interface using a precision razor blade exposed atomically smooth Au surfaces, with root-mean-square roughness about 0.3 nm 12. CTAC stabilized Au nanoparticles (80 nm diameter, $\sim{0.1}\;\rm{mg\cdot mL^{-1}}$ aqueous dispersion) were purchased from Micetech Co. Ltd. For NPoM assembly, $\sim 1\;\mu\rm{L}$ nanoparticle solution were drop-casted onto freshly stripped Au mirrors and incubated for 5 min under ambient conditions. Substrates were subsequently dried under nitrogen flow and immersed in ultrapure water with 10 s to remove excess CTAC on the sample surface, followed by secondary nitrogen drying. This protocol yielded NPoM structures with self-assembled $\simeq$ 1–2 nm CTAC spacer layers 12, forming well-defined plasmonic nanogap structures.
ASSOCIATED CONTENT
Notes
The authors declare no competing interests. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. A preprint version of this manuscript is available.32
Authors’ contributions
M. Kazemzadeh and B. Zhang contributed equally to this work. H. Hu conceived the idea and did electromagnetic simulations. M. Kazemzadeh trained the deep learning model and analyzed the data. B. Zhang, W. Jiang, and W. Chen performed the experimental investigation and collected data for the training. W. Chen, B. Zhang, T. He, H. Liu, Z. Jiang, Z. Hu, X. Dong, C. Sun, X. He, and H. Xu contributed to the establishment of the automatic collection optical setup. M. Kazemzadeh and H. Hu created the figures and wrote the original draft, reviewed and edited. All the authors discussed, reviewed, and edited the paper.
Data availability
All the source code, trained models, and the curated dataset are available at https://doi. org/10.5281/zenodo.19499234. All other data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
Funding Sources
This work was supported by the National Key Research and Development Program of China (Grant No. 2024YFA1409900) and the National Natural Science Foundation of China (Grant No. 62475071 and 52488301). M.K., and F.P. acknowledge funding by European Union’s Horizon Europe under grant agreement 101125498, project MINING - “Multifunctional nano-bio interfaces with deep brain regions”. M.K., and F. P. acknowledge funding from the European Union’s Horizon 2020 Research and Innovation Programme project DEEPER under grant agreement 101016787.
{suppinfo}
Please find a detailed explanation of the comparison of DF images and spectra of DL selections, UMAP unsupervised learning, SPARX architecture, homoscedastic SPARX model performance evaluation, generality test of SPARX on different plasmonic systems, and unsupervised learning on the nanoparticle shapes in the Supporting Information.